OpenStudy (anonymous):

Simplify 3 over x plus 2 over quantity x plus 2 all over 3 over quantity x plus 2 minus 2 over x Can anyone help me step by step?

OpenStudy (ranga):

Use the Equation button below the reply box to post your question using the mathematical editor so the problem is clearer.

OpenStudy (anonymous):

\[\frac{\frac{ 3 }{ x } +\frac{ 2 }{ x+2 } }{ \frac{ 3 }{ x+ 2 } - \frac{ 2 }{ x } }\]

OpenStudy (ranga):

\[\Large \frac{\frac{ 3 }{ x } +\frac{ 2 }{ x+2 } }{ \frac{ 3 }{ x+ 2 } - \frac{ 2 }{ x } } = \frac{ \frac{ 3(x+2)+2x }{ x(x+2) } }{ \frac{ 3x - 2(x+2) }{ x(x+2) } }\]

OpenStudy (ranga):

\[\Large = \frac{ \frac{ 5x+6 }{ x(x+2) } }{ \frac{ x - 4 }{ x(x+2) } } = \frac{ 5x+6 }{ x(x+2) } * \frac{ x(x+2) }{x-4} = \frac{ 5x + 6 }{ x-4 }\]

OpenStudy (anonymous):

Oh okay I see, the common factors cancel out! Thanks so much!!

OpenStudy (ranga):

You are welcome.